### v1

# The Authentication Gap: Higher Education’s Widespread Noncompliance with NIST Digital Identity Guidelines

Abstract
We examine the authentication practices of a diverse set of 101 colleges and universities in the United States and Canada to determine compliance with five standards in NIST Special Publication 800-63-3 Digital Identity Guidelines. We find widespread noncompliance with standards for password expiration, password composition rules, and knowledge-based authentication. Many institutions still require or recommend noncompliant practices despite years of expert advice and standards to the contrary. Furthermore, we observe that regional and liberal arts colleges have generally lower documented compliance rates than national and global universities, motivating further investment in authentication security at these institutions. These results are a wake-up call that expert cybersecurity recommendations are not sufficiently influencing the policies of higher education institutions, leaving the sector vulnerable to increasingly prevalent ransomware and other cyberattacks.
1 Introduction
Modern institutions are under constant threat of cyberattacks, including ransomware and phishing, that can completely cripple day-to-day operations, compromise sensitive data, and bear a hefty recovery cost [35]. Higher education institutions are especially vulnerable to these attacks as they process sensitive information and often lack resources to secure their infrastructure. In recent years, colleges and universities have been targeted by an increasing number of attacks [12]. A large portion of these attacks attempt to compromise authentication information of faculty, staff, or students in order to access and control sensitive data or resources.
To help mitigate potential attacks, the US National Institute of Standards and Technology (NIST) has published technical standards for digital authentication based on expert knowledge from the cybersecurity community [22]. Despite the availability and specificity of these standards and their updates over successive years, many institutions have been slow to adopt them in practice [18]. In 2022 and 2023, Lee et al. [27] and Hall et al. [24] found particularly low compliance with updated NIST authentication standards across the tech industry.
In this paper, we investigate compliance of higher education institutions with NIST Special Publication (SP) 800-63-3 Digital Identity Guidelines. We examine the authentication practices of a diverse set of 101 colleges and universities in the United States and Canada, including public, private, R1, liberal arts, and regional institutions. We review online documentation of these institutions’ authentication practices and check for descriptions of and compliance with five key standards in NIST SP 800-63-3: 1) password expiration, 2) password composition guidance, 3) password strength meters, 4) knowledge-based authentication (security questions), and 5) multi-factor authentication. The recommended best practices for each of these standards have changed since previous versions of the standard, making them ideal candidates for identifying noncompliance.
Among institutions with authentication policy details available online, we find an unfortunately large fraction are out of compliance with the NIST standard. Contrary to NIST’s guidelines, password expiration, knowledge-based authentication, and specific password composition rules are still widespread. The rates of noncompliance we observe are higher than in Hall et al. [24]’s 2023 study of websites across the tech industry, indicating that higher education is out of step with the broader online ecosystem on this issue. We do, however, observe widespread deployment of NIST-recommended multi-factor authentication (MFA) across the institutions we examine. This finding is in keeping with recent literature about engagement with and opinions of MFA at educational institutions [10, 4, 16].
Importantly, we note that noncompliance rates are not distributed evenly across institution categories. US liberal arts colleges, US national universities, and Canadian global universities have more online documentation of digital authentication policies than US regional colleges. Across institutions with online documentation, US regional colleges generally have lower compliance rates than US liberal arts colleges, which have generally lower compliance rates than US national and Canadian global universities. Although our data do not provide a causal explanation for this result, it is reasonable to assume that lower compliance rates of US regional colleges are related to their generally more limited resources on an overall and per-student basis.
Our results do provide direct evidence that investments to strengthen the US and Canadian infrastructure against cyberattacks would be well directed towards improving authentication security in the higher education sector, especially at less well-resourced institutions. While out of scope for this study, we hope that follow-up research will examine digital authentication security at additional institutional categories, including community colleges, historically Black colleges and universities (HBCUs), and Hispanic-serving institutions (HSIs).
2 Related Work
The rapid rise in computational capabilities and credential leaks that allow attackers to break a large number of passwords have forced researchers and practitioners to re-examine authentication policies, including forced password expiration, password composition rules, knowledge-based authentication, strength meters, and multi-factor authentication. Scholars have also examined the actual behavior of individuals and companies regarding these authentication practices; however, no prior study has compared authentication policies of higher education institutions to current industry standards as broadly as in this work.
Password Expiration.
Several prior efforts provided early evidence about the relative ineffectiveness of privacy expiration policies. In 2010, Zhang et al. [44] developed a framework that deduces new user passwords from old user passwords through a series of successive transformations. The framework was able to infer 41% of new passwords in an offline attack and 17% in an online attack. In 2014, Choong et al. [9] conducted a NIST study of US government employees’ password habits and found that when asked to create a new password, respondents tended to use less secure strategies, such as recycling old passwords or only making a minor change to an existing password.
In 2015, Farcasin and Chan-tin [17] surveyed university affiliates regarding pre-generated and expiring passwords. Respondents reported that a 120-day expiration time was too short, and the authors concluded that rapid expiration is untenable for most users, leading to password reuse and the creation of less secure new passwords. In 2018, Habib et al. [23] also surveyed users and found that regular password replacement usually led to similarly secure new passwords. They question the security gains of an expiration policy and recommend investing into alternative security measures.
In 2023, Gerlitz et al. [20] conducted a longitudinal study of employees of three German companies about the German Federal Office for Information Security’s removal of the password expiration requirement from their policy guidelines. While they reported a downward trend in requests to renew passwords, the investigation also revealed several factors that led to continued reliance of organizations on privacy expiration despite the federal recommendation against it. Several still, mistakenly, viewed the practice of periodically renewing passwords as beneficial to overall IT security. Several also kept password expiration while transitioning to MFA or were in industries with contradictory requirements, such as finance, which continue to mandate password expiration.
Knowledge-based Authentication.
Problems with knowledge-based authentication have been raised by the academic community for more than a decade. In 2009, Just and Aspinall [25] found that answers to security questions were typically low entropy (easier to crack) and users had trouble remembering their answers. In 2015, Bonneau et al. [6] examined personal knowledge questions at Google and found “a security level that is far lower than user-chosen passwords” with many users providing difficult-to-remember fake answers, concluding that “best practice should favor more reliable alternatives.”
Password Composition Rules.
Strict password composition rules are similarly problematic. While the idea of increasing password entropy through composition rules is good in theory, real users often make predictable choices that satisfy the composition rules but leave passwords vulnerable. In 2011, Komanduri et al. [26] found that composition rules mandating a mixture of cases, numbers, and symbols resulted in lower-entropy (worse) passwords than simply mandating longer passwords with no specific composition rules. In 2015, Ur et al. [38] found that many “weak passwords resulted from misconceptions, such as the belief that adding ‘!’ to the end of a password instantly makes it secure or that words that are difficult to spell are more secure than easy-to-spell words.” This misunderstanding is understandable given the widespread composition rule that passwords contain at least one symbol. Extensive composition rules also place a burden on users, resulting in less memorable passwords that are more likely saved in an insecure location or re-used for multiple accounts.
Multi-factor Authentication.
Adoption of multi-factor authentication (MFA) has accelerated in recent years, with online services from all sectors, including higher education [43] deploying optional or required MFA during login. In 2018, Colnago et al. [10] studied user opinions about MFA at Carnegie Mellon University. Users found MFA “annoying, but fairly easy to use, and believed it made their accounts more secure.” In 2022, Arnold et al. [4] similarly found that due to the “time sensitive nature of many tasks that required MFA, university students are likely to experience strong negative emotions towards MFA that drastically lower their perceptions of its utility and usability,” but that these emotions could be offset by an increased perception of security provided by MFA. These findings are corroborated by Dutson et al. [16], adding to a body of literature supporting widespread deployment of MFA at educational institutions.
Authentication Security in Higher Education.
A few studies have examined other aspects of authentication security in higher education. In 2022, Mayer et al. [28] studied why faculty, staff, and students at large educational institutions chose to use password managers or not. They found that perceived ease of use was the most important factor and recommended advocacy focusing on usability benefits. In 2023, Nisenoff et al. [31] found that many university accounts were vulnerable to credential-guessing attacks performed using cracked passwords from a data breach matched with email addresses.
Compliance with NIST Digital Identity Guidelines.
Relatively few studies have examined compliance with the 2017 NIST SP 800-63-3 Digital Identity Guidelines. In 2023, Hall et al. [24] analyzed over 100 websites across industries that “report the most breaches in the Verizon Data Breach Investigation Report.” They found a mixture of compliance and noncompliance, including “nearly all websites…avoiding the use of security questions and SMS-based 2FA” (a substantially higher compliance rate than the institutions in our study), but that “many websites (greater than 80 percent) still deem ‘P@ssw0rd’ an acceptable password.” In 2022, Lee et al. [27] “examined the [password] policies of 120 of the most popular websites” and “found that only 13% of websites followed all relevant best practices…75% of websites do not stop users from choosing the most common passwords…45% burden users by requiring specific character classes in their passwords for minimal security benefit” and “low [19%] adoption of password strength meters.” Our work is the first to examine the compliance of higher education institutions with NIST Digital Identity Guidelines specifically and at scale.
3 NIST SP 800-63-3 Background and
Research Questions
NIST Special Publication 800-63-3 Digital Identity Guidelines [22] was published in 2017 as an update to SP 800-63-2 (2013). The SP provides “an overview of identity frameworks; using authenticators, credentials, and assertions in a digital system; and a risk-based process to select assurance levels.” The SP was developed in collaboration with the community, during which it received over 1,400 comments in a draft period from 2016–2017. The digital authentication standards presented in NIST SP 800-63-3 are extensive and cover practices across the technical stack and of differing relevance to different types of organizations.
We selected five standards from SP 800-63-3 to investigate for this study. We chose these specific standards because they
  1. 1.
    Concern key aspects of the user-facing digital authentication process
  2. 2.
    Are widely accepted by the technical security community
  3. 3.
    Require relatively little expense to implement
  4. 4.
    Have changed or been introduced since previous versions of the SP
Factors 1–3 above mean that all institutions should be aware of and have implemented these standards since the publication of SP 800-63-3. Factor 4 means that these standards are useful checks for institutional responsiveness to technical guidelines, because practices following pre-2017 versions of the standard will be readily detectable.
The following subsections describe each of the standards we investigate in detail.
3.1 Password Expiration
In contrast to previous standards that online services should have their users change their passwords periodically, NIST SP 800-63B Section 5.1.1.2 paragraph 9 clearly states
Verifiers SHOULD NOT require memorized secrets to be changed arbitrarily (e.g., periodically). However, verifiers SHALL force a change if there is evidence of compromise of the authenticator. [21]
The previous wisdom held that the likelihood of password compromise increases over time and that regular password expiration is needed to reduce this risk. However, the NIST SP 800-63 FAQ provides the rationale for the new standard:
Users tend to choose weaker memorized secrets when they know that they will have to change them in the near future. When those changes do occur, they often select a secret that is similar to their old memorized secret by applying a set of common transformations such as increasing a number in the password. This practice provides a false sense of security if any of the previous secrets has been compromised since attackers can apply these same common transformations. But if there is evidence that the memorized secret has been compromised, such as by a breach of the verifier’s hashed password database or observed fraudulent activity, subscribers should be required to change their memorized secrets. However, this event-based change should occur rarely, so that they are less motivated to choose a weak secret with the knowledge that it will only be used for a limited period of time. [32]
We investigate whether higher education institutions require or recommend regular password expiration or cycling for their affiliates.
3.2 Knowledge-based Authentication
The NIST SP 800-63 FAQ states
Knowledge-based authentication (KBA), sometimes referred to as “security questions”, is no longer recognized as an acceptable authenticator by SP 800-63. This was formerly permitted and referred to as a “pre-registered knowledge token” in SP 800-63-2 and earlier editions. The ease with which an attacker can discover the answers to many KBA questions, and relatively small number of possible choices for many of them, cause KBA to have an unacceptably high risk of successful use by an attacker. [32]
As noted in this quote, the use of KBA for digital authentication was formerly permitted. The practice is still widespread. Most laypeople are familiar with the need to provide answers to questions such as “What was the model of your first car?” “What was your grandmother’s maiden name?” and “What is your favorite winter sport?” (as well as a wide variety of others) during account creation and authentication. The use of such questions is no longer acceptable as noted above and any such use would be considered out of compliance.
We investigate whether higher education institutions require or recommend KBA during account creation, login, and/or recovery.
3.3 Password Composition Rules
NIST SP 800-63B Section 5.1.1.2 states that
Verifiers SHOULD NOT impose other composition rules (e.g., requiring mixtures of different character types or prohibiting consecutively repeated characters) for memorized secrets. [21]
This is in contrast to previous standards that online services require specific composition rules during password creation. It is approaching common knowledge that all passwords “should” contain at least one symbol, number, and (often) mix of capital and lowercase letters. Most are familiar with password creation interfaces that enforce specific composition rules. However, these interfaces are all out of compliance with the current NIST standard.
NIST SP 800-63B Appendix A.3 provides the rationale for this change:
Research has shown, however, that users respond in very predictable ways to the requirements imposed by composition rules [Policies]. For example, a user that might have chosen “password” as their password would be relatively likely to choose “Password1” if required to include an uppercase letter and a number, or “Password1!” if a symbol is also required. [21]
We investigate whether higher education institutions require or recommend specific composition rules of this sort as a part of their password creation process.
3.4 Password Strength Meters
NIST SP 800-63B Section 5.1.1.2 states
Verifiers SHOULD offer guidance to the subscriber, such as a password-strength meter, to assist the user in choosing a strong memorized secret. This is particularly important…as it discourages trivial modification of listed (and likely very weak) memorized secrets [21]
NIST recognizes the need for authentication systems to help users create strong passwords and refers to strength meters as an example. Password strength meters are relatively common across online services, but have not received the same broad awareness as strict password composition rules. This is unfortunate, as password strength meters can be configured to account for a much wider range of weaknesses [13] that strict composition rules cannot prevent. Password strength meters are also more easily interpretable for users, often deploying colors and other interface elements to simplify guidance about a range of password vulnerabilities. However, password meters can be poorly configured, designed in a fashion that misinforms users, and may need to be supplemented with other forms of guidance. For example, some existing password meters may give the “green light” to weak passwords, such as “Jeremy1!” [18, 13].
We investigate whether higher education institutions provide password strength meters for their affiliates and require or recommend minimum strength values during password creation.
3.5 Multi-factor Authentication
According to NIST SP 800-63B Section 4.2, in order to achieve “authenticator assurance level 2” (“high confidence”) or stronger,
Authentication SHALL occur by the use of either a multi-factor authenticator or a combination of two single-factor authenticators. [21]
Multi-factor authentication (MFA) is well understood to be the best practice for digital authentication. Requiring two different forms of authentication, especially the combination of a memorized secret (e.g., a password) and a possession-based authenticator (i.e., “something you have”), makes account compromise significantly more challenging for would-be adversaries.
However, some authentication methods do not provide sufficient security to be deemed appropriate for use as a factor for MFA. NIST SP 800-63B Section 5.1.3.1 states that
[Authentication] methods that do not prove possession of a specific device, such as voice-over-IP (VOIP) or email, SHALL NOT be used for out-of-band authentication. [21]
We therefore investigate whether or not the institutions examined in this paper 1) recommend and/or require MFA and 2) allow email as an authentication factor.
4 Methods
This section presents our methods, including which higher education institutions we chose to investigate and why, how we systematically reviewed the publicly available online policies of these institutions for details about the digital authentication practices of interest (Section 3), our protocol for contacting institutional representatives to confirm or update the collected data, and some limitations that may arise from the nuances of our method.
4.1 Institution Selection
There are over 5500 higher education institutions in the United States [30] and over 400 in Canada [11]. Due to resource constraints that made investigation of the entire higher education sector infeasible, we selected a limited set of institutions as the focus of this study. It has been well observed that the actions of a few well-regarded institutions often have an outsize influence on behavior across the higher education space. We therefore adopted a selection process that prioritized well-regarded institutions from several broad categories.
Specifically, we selected 101 institutions from the U.S. News and World Report lists of “top” institutions from several categories as of September 2023. We chose the top 20 US national universities [39], top 20 US liberal arts colleges [40], top 15 Canadian global universities [41], and a selection of US regional colleges consisting of the top 10 in each of the North, South, Midwest, and West regions [42]. Since multiple institutions may tie for a ranking on this list, this selection process resulted in 101 total institutions. The complete list of institutions is provided in Table 1.
US National Universities US Liberal Arts Colleges US North and South Regional Colleges US Midwest and West Regional Colleges Canadian Global Universities
Princeton University Williams College United States Coast Guard Academy Illinois Wesleyan University University of Toronto
Massachusetts Institute of Technology Amherst College Cooper Union for the Advancement of Science and Art Ohio Northern University University of British Columbia
Harvard University United States Naval Academy United States Merchant Marine Academy College of the Ozarks McGill University
Stanford University Pomona College Grove City College Taylor University University of Alberta
Yale University Swarthmore College Maine Maritime Academy Simpson College McMaster University
University of Pennsylvania Wellesley College Pennsylvania College of Technology Cottey College Universite de Montreal
California Institute of Technology United States Air Force Academy Elmira College Alma College University of Calgary
Duke University United States Military Academy at West Point Alfred State College–SUNY Benedictine College University of Waterloo
Brown University Bowdoin College College of Mount St. Vincent William Jewell College University of Ottawa
Johns Hopkins University Carleton College SUNY College of Technology at Canton Hiram College Western University
Northwestern University Barnard College University of Maine at Farmington Lake Superior State University Dalhousie University
Columbia University Claremont McKenna College High Point University Embry-Riddle Aeronautical University–Prescott Simon Fraser University
Cornell University Grinnell College Florida Polytechnic University California State University–Maritime Academy University of Victoria
University of Chicago Middlebury College Beacon College Carroll College University of Manitoba
University of California, Berkeley Wesleyan University Flagler College Criswell College Laval University
University of California, Los Angeles Davidson College Wesleyan College Oregon Institute of Technology
Rice University Hamilton College Catawba College Fashion Institute of Design & Merchandising
Dartmouth College Harvey Mudd College University of the Ozarks College of Idaho
Vanderbilt University Smith College Spring Hill College Brigham Young University–Hawaii
University of Notre Dame Vassar College Huntingdon College Brigham Young University–Idaho
Colgate University Barton College University of Antelope Valley
Haverford College Newberry College
Washington and Lee University
Table 1: Higher education institutions included in this study. These institutions were selected from the 2023 U.S. News and World Report lists of top institutions in their respective categories [39, 40, 42, 41].
We recognize that this particular list of institutions is not without limitations, which we discuss in greater detail in Section 4.5. We strongly encourage follow-up studies investigating the digital authentication policies of different institutions, e.g., public flagship universities, historically Black colleges and universities (HBCUs), Hispanic-serving institutions (HSIs), and community colleges.
4.2 Data Collection
The data collection for this study followed a straightforward protocol for all selected institutions. Data collection took place between September 2023 and August 2024.
4.2.1 Data Format
We first identified the specific digital authentication standards from NIST SP 800-63-3 listed in Section 3 as the focus of our study. We then created a spreadsheet to support standardized data collection across standards and institutions. For each standard, we sought to answer the following question for each institution and code the answer as one of a limited set of data codes:
  1. 1.
    Does the practice associated with this standard appear in the publicly available online policies of the institution?
    1. (a)
      “No Policy Found": There is no information about this specific practice
    2. (b)
      “In Policy”: There is information about this specific practice
If the answer to question 1 was “in policy,” we then sought to answer the following question:
  1. 2.
    What is the institutional policy regarding this practice?
    1. (a)
      “Required”: The practice is required for all affiliates of the institution
    2. (b)
      “Required for Specific Affiliates”: Some affiliates of the institution are required to follow the practice. In this case, we also recorded which specific affiliates are subject to this requirement.
    3. (c)
      “Recommended”: Affiliates of the institution are recommended to follow the practice, but adherence is not enforced.
    4. (d)
      “Discouraged”: Affiliates of the institution are discouraged from following the practice, but prevention is not enforced.
    5. (e)
      “Disallowed”: Affiliates of the institution are not permitted to follow the practice.
We chose the answer codes for both questions after a trial data collection with three institutions. The codes covered all cases we found throughout the collection process. In addition to this consistently formatted data, we also took open-ended notes on any particularly interesting policies we discovered.
4.2.2 Data Sources
We sought to answer questions 1 and 2 above based on any publicly available online policies posted on official institutional websites. We used several methods to locate these policies, including (but not limited to):
  1. 1.
    Search engines (e.g., Google) with site-specific search queries (“site:[institution].edu”)
  2. 2.
    Search bars on institutional websites
  3. 3.
    Manual navigation of institutional websites (e.g., looking for links to policies or documents about digital authentication)
This process was intentionally open-ended, reflecting the variety of institutional website structures and information posting practices. We found relevant policies most frequently in the following documents:
  1. 1.
    An “information security policy,” “cybersecurity policy,” or “IT policy” often posted by the IT department of the institution
  2. 2.
    IT “support pages” (also referred to as “guides,” “help desk,” “knowledge base,” or similar) posted by the IT department of the institution
  3. 3.
    A “privacy policy” or “privacy notice” often posted on the homepage of the institution
At least one author and a trained research assistant manually reviewed each of the relevant policies, recording answer codes for each standard.
4.3 Data Confirmation
Publicly available online information posted by institutions is not always up to date with actual practices. We therefore reached out to representatives at all studied institutions according to the below protocol for confirmation and/or updates to the information we gathered.
We first identified a contact email address for the information technology department at the institution. For most institutions, this was the email to an IT help desk or equivalent (e.g., “ithelp@” or “servicedesk@”). Some institutions had specific email addresses for security-related contacts, which we used if available (e.g. “information.security@”). Finally, a few institutional web sites directed queries directly to a specific IT staff member.
We then emailed all institutions with a customized email including an introduction to our study, a summary of the information we had already collected about the institution from their publicly available policies, and a request to complete a Google Form with any corrections, clarifications, or completions of missing data. We indicated that such responses would be appreciated and would ensure that the institution was not misrepresented in our paper.
Out of the 101 institutions we included in this study, only 13 responded to our outreach. Of these, 9 said that they could not provide any information for our study. Typical responses were akin to “Unfortunately we are not able to participate in this project” and “ITS will not respond to requests outside of their institution.” Only 4 institutions actually provided useful information by completing the Google form.
Most of these form responses corroborated the data we had already collected from the institutions’ online policies. In the case of discrepancies, we report the data from the online policies (rather than the form responses). This paper therefore provides a snapshot of publicly available online policies across the sector, unbiased by form responses from a small number of institutional representatives (who are incentivized to provide socially correct descriptions of their security practices).
4.4 Research Ethics
The Colgate University Institutional Review Board determined that this study was exempt. We only report data from publicly available sources posted online by the institutions we study. Although these data are not private, we anonymize our results, refraining from naming-and-shaming any particular institutions.
As previously discussed, the limited amount of data provided directly by the institutions for data confirmation purposes was not included in our analysis. We do not collect or report any information about individual affiliates at any institution.
4.5 Limitations
The methods used in this paper have some limitations that should be acknowledged to understand the scope of the results.
First, it is essential to remember that this study examines institutional policies, and that these policies may or may not represent actual, technologically-enforced institutional practices. For example, a policy might say that multi-factor authentication is required for all institution affiliates, but some affiliates might, in practice, be able to disable MFA via their internal profile settings. Similarly, a policy might state that yearly password expiration is required for all affiliates, but the current IT leadership might not actually enforce regular password cycling. Nevertheless, we expect posted policies to generally align with actual practices, especially since posted policies are often primarily intended as guides for affiliates (e.g., on IT “help” websites). The policies signal the intentions and values of the institution and influence the actions of affiliates and peer institutions. On a practical level, determining actual technically-enforced practices would require institutional affiliation or direct input from institutional representatives, which would be challenging given the low response rate to our data confirmation requests and hostility to discussing cybersecurity practices in most of the responses we did receive.
Second, the 101 institutions we studied represent only a fraction of the higher education sector. As noted above, we encourage follow-up studies examining digital authentication practices and other forms of cybersecurity compliance across other countries and institutional categories. The institutions in this paper are at the top end of prestige in their respective categories, which must be remembered when interpreting the results. These institutions, in contrast to less prestigious peers, are more likely to be well-resourced, making them more likely to have sufficient staff with expertise and availability to ensure the institution’s digital authentication (and other cybersecurity) practices follow accepted guidelines. The fact that we find many cases of noncompliance with NIST standards among these prestigious institutions’ online policies suggests that the overall rate of noncompliance is substantially higher across the higher education space.
Third, this study reflects a snapshot in time, with data collected between September 2023 and August 2024. Follow-up longitudinal studies are necessary to see whether the trajectory of institutional policies and practices is trending toward compliance with NIST standards.
5 Results
This section presents the results from our investigation of the publicly available online policies of 101 institutions of higher learning in the United States and Canada.
For each authentication practice described in Section 3, we first report what fraction of institutions in each category refer to the authentication practice in their online policies. We then focus on the subset of institutions with such information and report the prevalence of specific practices. We note which practices are in or out of compliance with NIST SP 800-63-3 and provide anecdotal examples.
Figure 1: Online publicly available policies of 101 higher education institutions regarding standards from NIST SP 800-63-3. Policies in green align with the standards. Policies in red are noncompliant.
5.1 Password Expiration
We found that requiring or recommending password expiration/cycling is still widespread across higher education (Figure 1). 47% (47) of all investigated institutions referred to password expiration in their online policies. Of these, 47% (22) required regular password expiration for all affiliates, 17% (8) required regular password expiration for specific affiliates (faculty/staff, administration, or “HIPAA affectees”), and 28% (13) recommend (but did not require) regular password expiration for all affiliates. These policies are in direct violation of the NIST 800-63-3 standard that password expiration should only occur in response to a known breach. Only 11% (5) of institutions that referred to password expiration in their online policies did so to explicitly disallow or discourage the practice.
The institutions that require or recommend password expiration for some or all affiliates do so for a wide range of expiration frequencies (Table 2). One year was the most common password expiration frequency we observed, but specific frequencies ranged from expiration every 1 month to every 400 days, with one institution recommending password expiration “often,” one recommending “periodically,” and several not listing a frequency in their publicly available policy.
Password Expiration Frequency Count
1, 3, or 6 months 1
60 days 2
90 days 3
3 months 2
3-6 months 1
120 days 4
126 days 1
180 days 4
Once every semester 1
6 months 4
365 days 1
1 year 12
400 days 1
Periodically 1
Often 1
Table 2: Password expiration frequencies for institutions that require or recommend regular password cycling.
Some of these institutions purposefully disregard the NIST standard. One university has a Q&A on “Why Are We Implementing A Password Change Initiative When It Is No Longer Considered Best-Practice?” that states
Changing passwords regularly and implementing a password expiry date helps to limit the use of compromised accounts by attackers for malicious activities. In an effort to provide better account management while adhering to the spirit of best practice guidance, a reset interval that is longer than “90 days” but shorter then “never” is being put into practice.
A smaller percentage of US regional colleges (37%) and Canadian universities (40%) referred to password expiration in their policies than institutions in other categories ($\geq$60%). However, across institutions with policies referring to password expiration, a greater fraction of US regional colleges required expiration for all affiliates (67%) versus institutions in other categories ($\leq$50%).
5.2 Knowledge-based Authentication
The use of security questions (e.g. “What is the name of your first pet?”) during the authentication process remains alarmingly widespread across higher education (Figure 1). 45% (45) of all investigated institutions referred to the use of security questions during account creation, login, and/or recovery in their online policies. Of these, 62% (28) required security questions for all affiliates, 25% (11) required security questions for specific affiliates, and 11% (5) recommended (but did not require) security questions for all affiliates. All such policies are in violation of the NIST SP 800-63-3 standard that security questions are no longer recommended for any digital authentication process due to their inherent insecurity. Only a single institution discouraged the use of security questions.
Overall, a smaller percentage of US regional colleges (30%) referred to security questions in their policies than institutions in other categories (>>>52%). Across institutions with policies referring to security questions, a smaller percentage of US universities actually require them for all affiliates (33%) as compared to institutions in other categories (>>>69%).
The institutions that use security questions typically do not provide any guidance about which questions or what types of answers are likely to be more secure. One institution described transitioning from requiring one security question to requiring three security questions. While this approach may be slightly more secure, it is still out of compliance with the NIST standard.
5.3 Password Composition Rules
Despite the NIST SP 600-83-3 standard that institutions should not impose password composition rules because they “do not significantly improve the security of selected passwords” [32], the presentation and enforcement of password composition rules appears widespread across higher education (Figure 1). 73% (74) of all investigated institutions referred to password composition rules in their online policies. Of these, 77% (57) require that affiliates meet minimum password composition rules during the password creation process. Another 22% (17) recommend (but do not require) that affiliates meet password composition rules.
Similar to previous cases, a smaller percentage of US regional colleges (58%) referred to password composition rules than institutions in other categories ($\geq$80%). Across institutions with policies referring to password composition rules, greater than 80% of US institutions across categories enforce compliance with these rules during password creation. A much smaller percentage of Canadian universities referring to password composition rules explicitly require compliance (41%). Figure 2 shows a representative example of password composition rules from a US regional college.
Figure 2: Example password composition rules from a US regional college.
5.4 Password Strength Meters
Rather than specific password composition rules, NIST SP 800-63-3 recommends that institutions provide graphical password “strength meters” during the password creation process. Unfortunately, only 6% (6) of all investigated higher education institutions referred to password strength meters in their online policies (Figure 1). This low fraction is effectively uniform across institution categories. Of this fraction, one US national university and one US regional college require that affiliates meet some minimum strength requirement on a meter during password creation (e.g., a rating of “strong” on the meter shown in Figure 3 from a US regional college).
Figure 3: Example strength meter from a US regional college that requires a rating of “strong” during password creation.
5.5 Multi-factor Authentication
The use of multi-factor authentication (MFA) is widespread in higher education (Figure 1). 76% (77) of all investigated institutions referred to MFA in their online policies. Of these, 73% (56) required MFA for all affiliates, in compliance with NIST SP 800-63-3. A further 9% (7) required MFA for specific affiliates, and 16% (12) recommend (but did not require) MFA for all affiliates. The institutions that only require MFA for certain affiliates do so for employees only (i.e., not students) or only for IT staff.
A considerably smaller percentage of US regional colleges referred to MFA in their online policies than institutions in other categories (51% vs >>>90%). Of the institutions with policies referring to MFA, US national universities have the highest requirement rate for all affiliates (84%), followed by US liberal arts colleges (76%), Canadian global universities (71%), and US regional colleges (64%).
NIST SP 800-63-3 further specifies that email should not be considered as a valid second factor for multi-factor authentication. 91% (92) of the institutions did not provide any information about the use of email as a second factor in their online policies. Fortunately, many of these institutions use a third-party service for MFA, such as Duo, that does not allow email as a second factor [15]. However, 7% (7) of investigated institutions do indicate that email is permitted as a second factor (Figure 1).
5.6 References to NIST Standards
37% (37) of investigated institutions referenced NIST standards related to authentication security in their online policies (Figure 4). This was most common for Canadian global universities (67%) and US national universities (55%) and least common for US regional colleges (19%).
In some cases, these references acknowledge some noncompliance. For example,
[Anonymized] administration recognizes that fully implementing all controls within the NIST Standards is not possible due to institution limitations and resource constraints. (US regional college)
Others explicitly claim that the institution will adhere to the standard. For example,
[Anonymized] University will: Follow the National Institute of Standards and Technology Framework for Improving Critical Infrastructure Cybersecurity (NIST Cybersecurity Framework). (US national university)
While many others just reference NIST as a guiding resource. For example,
This policy framework consists of eighteen (18) separate policy statements, with supporting Standards documents, based on guidance provided by the National Institute of Standards and Technology. (US liberal arts college)
Figure 4: Whether 101 higher education institutions reference NIST standards related to authentication security in their online policies.
6 Discussion
The results of this study reveal an alarmingly large gap between many higher education institutions’ authentication policies and the NIST Digital Identity Guidelines. In this section, we discuss the implications of these results.
Cyberattack Vulnerability.
Our results are particularly worrisome given the increasing prevalence of cyberattacks on educational institutions. The increased use of remote learning and e-learning platforms has increased the threat surface of higher education [5]. According to Alexei and Alexei [2], the threat of “DoS / DDoS attacks, cross-site scripting, spoofing, unauthorized data access and infection with malicious programs, [and] also the theft of personal data has increased dramatically.” Ransomware [36, 37, 1] and phishing [7, 14, 3] have also become more common, with ransomware attack prevalence increasing over 70% in 2023 [12] and recovery costs averaging around \$1 million [35].
The continued use of insecure authentication practices makes institutions even more susceptible to cyberattacks. The NIST compliance rates we observe for higher education institutions is considerably lower than that observed by Hall et al. [24] across a swathe of industries identified as prone to data breaches. This should be a wake up call to higher education IT departments, administration, and all stakeholders about the still limited adoption of the NIST standards, all of which reflect well known best practices. As noted in Section 3, all of these standards are well within the expertise of IT departments to implement, especially given the availability of relevant code libraries and third-party services.
Differences Across Institutional Categories.
We notice that across the NIST standards we investigate, US regional colleges provide less online documentation than institutions in the other categories (more on this below). However, when considering the institutions that do provide documentation, US regional colleges have compliance rates below US liberal arts colleges, which themselves have compliance rates below US national and Canadian global universities. Of the institutions with policies about password expiration, US regional colleges were also the only category with greater than 50% requiring regular password expiration for all affiliates. US regional colleges were also the only category we examined with fewer than 50% requiring MFA overall.
While our results do not provide a causal explanation for this correlation, regional colleges are typically less well-resourced than their top-ranked university or liberal arts counterparts. This could lead to fewer or more overburdened IT staff, less money for third-party MFA solutions, etc. We expect that other categories of traditionally lower-resourced institutions (e.g., community colleges, HSIs, and HBCUs) also have low rates of compliance and encourage additional research for verification.
This finding is important because it provides direct guidance to focus government and industry cybersecurity initiatives. Meaningful gains in cyberattack threat reduction could be made with investment in improving digital authentication security for lower-resourced higher education institutions already less able to afford expensive attack recovery processes. IT conferences, peer workshops, and other fora should be leveraged to help bring all institutions in line with best practices as quickly as possible.
Limited Policy Availability.
The availability of online documentation about authentication practices varied widely across the standards we examined. Very few institutions had posted any information about their use or disuse of password strength meters or email as a second factor for MFA. Approximately 45% of institutions included something about password expiration or knowledge-based authentication in their online documentation. Over 75% of institutions posted about password composition rules or MFA.
The implications of this variety of posted policies are difficult to interpret. For example, consider an institution that does not have an online policy about MFA. It likely means that the institution does not support MFA and therefore does not need to have a MFA policy. However, it is possible that the university does support MFA, but the instructions for setting up and using the service are only available to authenticated affiliates. As an alternative example, consider an institution that does not have an online policy about password composition rules. In this case, one might assume that the institution does not enforce such rules; however, password composition rules are so widespread that it seems unwise make an inference either way. For example, such an institution might have a password symbol/number/case requirement that only appears when already-authenticated users attempt to change their passwords.
Nevertheless, we did observe one consistent trend in the availability of online policies: Across all standards we examined, a greater percentage of US regional colleges had no publicly available policies than institutions in any other category. This is notable, because digital authentication policies ideally should be publicly available online, periodically updated to reflect current practices, and standardized so that disclosures communicate information sufficient to inform relevant stakeholders. Universal compliant disclosure of such policies would serve as a deterrent to cyberattacks and a motivator for other organizations to follow similar practices. It would also serve an educational purpose, teaching affiliates about the authentication practices they should understand and expect in other aspects of their online lives. We hope that this research incentivizes more institutions to bring their digital authentication practices into compliance with known best practices and provide public notification that they have done so.
To simplify this process, we advocate for the standardization of public-facing authentication policy notices, e.g., via a free template. This would make it simple for institutions to 1) see what is necessary to bring their practices into compliance in order to use the template, and 2) quickly create a public-facing document describing these practices with a widely-recognized, user-friendly, and ideally machine readable format. Such a template should allow easy customization for details unique to an institution (e.g., what specific third-party MFA provider is in use) within constraints ensuring that the resulting notice still describes compliant practices. This would be analogous to the standardization of privacy policies and terms and conditions documents, but hopefully with a greater emphasis on readability. The specifics of this standardization would need to be developed in conversation with higher education institutions, but preliminary versions could be proposed and tested in future academic cybersecurity and human-computer interaction research.
Data Confirmation Difficulty.
As noted in Section 4, online documentation does not always reflect institutional practice. However, only 4 institutions replied to our outreach with meaningful data confirmation or updates. As we discussed, we do not believe that this limits the results of this research. In most cases, online documentation will match practice. IT departments know that community members look to the documentation when they have questions or issues with authentication, so there is an incentive for it to be accurate.
Furthermore, it is useful to evaluate the publicly available policies of institutions whether or not they exactly align with current practice. These policies signal an institution’s values and priorities. If an institution leaves an outdated, noncompliant policy posted online for affiliates and the general public to view, it signals to other institutions that such practices are still acceptable. This can limit the adoption of new best practices and extend existing vulnerabilities.
Nevertheless, the difficulty we faced confirming actual digital authentication practices reflects the difficulty of research in the higher education security and privacy space. Anecdotal conversations with other higher education cybersecurity and privacy researchers suggest that the difficulty of obtaining data is severely limiting the field. Institutions are hesitant to share details of cybersecurity practices with outsiders even if the secrecy of the practices is irrelevant to actual institutional security (c.f. Kerckhoffs’s principle [33]). One IT security director even contacted the authors to apologise that while they would have liked to share more details about their institution’s practices, they had been forbidden to do so by those higher in administration.
We hope that the resistance towards sharing cybersecurity data with academic researchers abates with continued advocacy. The very same institutions that balk at contributing to academic research vetted by institutional review boards readily contract with private education technology platforms having known privacy and security vulnerabilities [34]. Given the breadth of outdated authentication policies we observe in this study, more effort is needed to incentivize institutions to update their practices and better protect their affiliates and operations.
Governance Considerations: Incentives and Knowledge.
Our study of organizational compliance with NIST standards could be framed and understood strictly in terms of incentives, resources, and priorities. From this perspective, the NIST standards constitute non-binding, technical guidance about authentication and digital identity that informs cost-benefit calculations, risk assessment/management, procurement, and other organizational decision making processes. Higher education organizations are complex, and various departments and personnel may be responsible for taking NIST standards into account. Resource constraints play a substantial role in cybersecurity, and as we noted above, variations in resources across institutions likely impacts (non)compliance with NIST standards. To increase the rate of compliance and consequently improve cybersecurity in higher education, society might need to re-engineer the relevant incentives, resources, and priorities. This might entail increasing public funding directed towards better cybersecurity practices (e.g., to support MFA), improving disclosure of policies and practices along the lines suggested previously (which might impact accountability to different stakeholders), and adjusting incentives and priorities through external pressure (e.g., insurance, markets, even legal reform).
However, this perspective does not provide a fully satisfactory explanation for the high rate of noncompliance with NIST standards that are quite easy and low cost to implement. Password expiration is the most obvious example. Refraining from requiring or recommending password replacement, except when a breach has occurred, is not difficult or expensive. Compliance is easy and should be widespread, yet our findings suggest it is not. We therefore believe the reason is not solely rooted in misaligned incentives, lack of resources, and alternative priorities. Instead, we think that this area is subject to a knowledge commons problem [19] concerning the diffusion of evolving expert knowledge, i.e., how such evolving knowledge translates (or not) into professional practice and system design.
The evolution of expert knowledge about authentication security over the course of decades has led to multiple 180-degree shifts in important NIST standards. The 2004 NIST SP 800-63 Electronic Authentication Guideline [8] advised users to protect and secure their accounts with complex passwords that comprise random characters, capital letters, and numbers, and to change their passwords regularly [29]. Federal agencies, corporations, and universities largely followed the advice by enforcing password complexity (e.g., mixtures of numbers, letters, and symbols), length, and/or regular password expiration. But expert knowledge on authentication systems changed dramatically as the field broadened beyond computational cybersecurity to include human factors, usability, and other knowledge sources [18]. Recognizing flaws and limitations of the 2004 publication, NIST published updates in 2011 (“800-63-1”), 2013 (“800-63-2”), and 2017 (“800-63-3”) [22]. In these successive updates, NIST established new guidelines for authentication to replace widespread practices that undermined security. One 180-degree revision, for example, was for organizations to no longer require periodic password expiration; instead, users should only change their passwords if there is evidence of compromise [21].
Frischmann and Johnson [18] suggest that our observed noncompliance with these updated recommendations may be attributable to knowledge commons dilemmas that inhibit diffusion of evolving expert security knowledge, initially to intermediate practitioners/professionals, such as software vendors and IT departments, and ultimately to laypeople. From this perspective, our study suggests that knowledge of changes in NIST standards may not have effectively diffused to many higher education institutions. If noncompliance is a consequence of delayed diffusion of knowledge, it may work itself out over time as more individuals gradually become aware of the updated wisdom. Alternatively, new forms of engagement may be necessary to facilitate the timely diffusion of expert knowledge reflected in NIST standards to those actually responsible for setting and implementing authentication policies and practices. This could involve continuing education and training courses at college and university IT departments, automated identification and notification of noncompliance by researchers or watchdog groups, or other to-be-developed methods of disseminating expert knowledge.
7 Conclusion
We examined online policies of 101 higher education institutions in the US and Canada to measure compliance with NIST SP 800-63-3 Digital Identity Guidelines. We focused on five standards that reflect changes from prior wisdom and impact all institutional account holders. We found broad deployment of multi-factor authentication, but widespread noncompliant use of password expiration, knowledge-based authentication, and password composition rules. These issues leave institutions vulnerable to cyberattacks, and less well-resourced institutions seem to have greater rates of noncompliance.
These results serve as a wake-up call that best practices that are well-understood by experts in the cybersecurity community have not sufficiently influenced the policies of higher education institutions. More investment and outreach in this area is needed, especially as improved authentication practices will harden institutions against the rising prevalence of ransomware and other cyberattacks.
Acknowledgments
We thank Julia Tuck for her contributions. This research was supported by the Colgate University Faculty Research Council.